The empirical formula of a compound is CH. At 200 degrees Celsius, 0.242 of this compound occupies 97.2 mL at a pressure of 0.740 atm. What is the molecular formula of the compound
1
Expert's answer
2019-10-22T05:43:33-0400

M=mRT/PV=0.242*8.314*(200+273.15)/(0.74*101.325*0.0972)=130

(number of CH)=M/M(CH)=130/(12+1)=10

Answer: C10H10.